OpenStudy (ilovecake):

What is the value of the expression z -4/-8 when z = –28?

OpenStudy (anonymous):

Why is yours yellow?

OpenStudy (campbell_st):

so replace z in the expression with -28 and you get is the expression 1. \[\frac{z - 4}{-8}\] or 2. \[z - \frac{4}{-8}\]

OpenStudy (ilovecake):

This is a QH question..

OpenStudy (dan815):

Hello, look at campbells question is it the first or the 2nd one

OpenStudy (ilovecake):

It is number 1.

OpenStudy (campbell_st):

so replace z with - 28 \[\frac{-28 - 4}{-8}\] simplify the numerator... then divide the answer by the denominator... hope it helps

OpenStudy (anonymous):

Find the volume of the square pyramid to the nearest cubic unit. edge 11 ft; height 11 ft what would this be

OpenStudy (ilovecake):

I got 4 as the answer.

OpenStudy (campbell_st):

correct